# GAD-Generative Learning for HD Map-Free Autonomous Driving

### v2

@@ A.11 Road Test
The route then follows a right-turning at the next intersection and enters Huocheng Road, a narrow one-lane-per-direction road where trucks from the opposite direction sometimes encroaching on the ego lane, necessitating slight steering to avoid collisions. Proceeding through the next intersection, the route goes straight and executes an unprotected left turn to cross a bustling transportation route frequently used by trucks. Finally, the route circles back to Yecheng Road and returns to destination No.1688. The total length of the test route is 3.0 kilometers long.
A.12 Dateset
We collected the record files in 15 hours, which captured the perception output, the ADV trajectory, and route info. All the records are post-processed to obtain 200k clips in total. By default, we take the 1.5s of past context and predict/plan the future 3.0s contexts, corresponding to 15 frames in the past and 30 frames in the future. We partition the dataset into 16w for training and 4w for testing.
@@ A.13 Training details
In the first stage of our proposed two-stage training model, we regress a grid map by a cnn encoder-decoder backbone, which represents the occupation that vehicles may distribute in the future. Specifically, cnn-encoder embeds the scene image into feature map with 12 convolution layers, and then 4 deconvolution layers are used to decode feature map to the output map.
In the second stage of obstacle trajectory prediction, we reuse the CNN encoding layer from the first stage and initialize the encoder parameters with the last training step. As for obstacle trajectory prediction, we use an LSTM layer to replace the cnn decoder. In order to concentrate more on obstacles, we crop the encoding feature map centered agents, as the local scene feature of obstacles. The cropping direction aligns with the orientation of the obstacles. The local scene feature map will be embedding as a local scene vector through a Multi-Layer Perceptron layer and a max-pooling layer. Then, the local scene vector will be sent to the LSTM decoder, concatenated with agent’s history vectorized embedding vector.

### v3

@@ A.15 Road Test
The route then follows a right-turning at the next intersection and enters Huocheng Road, a narrow one-lane-per-direction road where trucks from the opposite direction sometimes encroaching on the ego lane, necessitating slight steering to avoid collisions. Proceeding through the next intersection, the route goes straight and executes an unprotected left turn to cross a bustling transportation route frequently used by trucks. Finally, the route circles back to Yecheng Road and returns to destination No.1688. The total length of the test route is about 3.0 kilometers long.
The video for road test is available at https://youtu.be/DYPB5lCRvn0
A.16 Dataset
@@ A.17 Training details
In the first stage of our proposed two-stage training model, we regress a grid map and cost volume by a CNN encoder-decoder backbone, which represents the occupation that vehicles may distribute in the future. Specifically, CNN-encoder embeds the scene image into feature map with 12 convolution layers, and then 4 deconvolution layers are used to decode feature map to the output map.
In the second stage of obstacle trajectory prediction, we reuse the CNN encoding layer from the first stage and initialize the encoder parameters with the last training step. As for obstacle trajectory prediction, we use an LSTM layer to replace the CNN-decoder. In order to concentrate more on obstacles, we crop the encoding feature map centered agents, as the local scene feature of obstacles. The cropping direction aligns with the orientation of the obstacles. The local scene feature map will be embedding as a local scene vector through a Multi-Layer Perceptron layer and a max-pooling layer. Then, the local scene vector will be sent to the LSTM decoder, concatenated with agent’s history vectorized embedding vector.
The training methodology for the GAN-based planner follow similar two-stage procedural, undergoing fine-tuning on the pre-trained imitation planner to stabilize and accelerate the training process.
A.18 Qualitative Results
We show typical qualitative results in Fig. 12 - Fig. 16 to illustrate that our work ensures the ADV can navigate through diverse traffic scenarios to reach the destination safely. Note that (a) shows the traffic scenario, and the blue points are the final trajectory selected by our model. (b) is the visualization of the cost volume which explains the areas with higher probabilities to travel. (c) is the visualization of the grid map to illustrate the future occupancies of obstacles. It should be noted that, figure (b) shows the superposition of our predictions for the occupancy at the next 1s, 2s, and 3s. The red area represents the highest probability of occupancy for the next 1s for the ego vehicle. The green and blue areas represent the highest probability of occupancy for the next 2s and 3s, respectively. Figure (c) illustrates the superposition of occupancy at the current moment, in 1s, and in 2s. The red region denotes the current position of the obstacle, while the green indicates the areas that the obstacle may occupy at 1s, and the blue represents the areas the obstacle may occupy at 2s.
Figure 12: Qualitative results: as the vehicle in the right lane accelerates and merges into the ego lane, the ADV opts to decelerate in order to prevent potential rear-end collisions.
Figure 13: Qualitative results: the ADV is transitioning to the rightmost lane, while also considering the option of using the middle lane if the rightmost lane is occupied.
Figure 14: Qualitative results: the ADV is maneuvering out of the traffic congestion by overtaking the preceding vehicle at a low speed in order to mitigate potential hazards.
Figure 15: Qualitative results: the ADV is performing a right turn at a decreased speed to prioritize safety.
Figure 16: Qualitative results: the ADV is making a left turn at an unprotected intersection. As a vehicle from the oncoming direction approaches, the ADV reduces its speed.